OpenStudy (anonymous):

Find the magnitude and direction of the resultant vector A+B, A mag:8m 270deg, B mag:15 60deg

OpenStudy (anonymous):

Do you know how to decompose into horizontal and vertical components?

OpenStudy (anonymous):

i believe so...for A I got x=0, y=8, z=0 for B I got x=(15/2), y=15[\sqrt{3}\]/2, z=0

OpenStudy (anonymous):

Yes, this is two-dimensional, so you can disregard the z's. Careful for that first one, though. If the angle is 270º, y= -8.

OpenStudy (anonymous):

i have a picture of my work if that would help

OpenStudy (anonymous):

thought so with the -8

OpenStudy (anonymous):

Yeah, just change that +8 to -8 and you should have it. This is what I got: R = 9.01m at 33.69º

OpenStudy (anonymous):

Thank you. I will try it again.